# Climbing the Ladder of Interpretability with Counterfactual Concept Bottleneck Models

### v1

@@ B.2 Evaluation Metrics
In our analysis we use the following metrics. What? (prediction generalization): we compute the Area Under the Receiver Operating Characteristic Curve [Hand and Till, 2001] for concepts and tasks (ROC AUC ($\uparrow$). Why? (explanation quality): we compute the impact of each concept on the downstream task, inspecting the task predictor $f$ [Yuksekgonul et al., 2022]. For each class, we look at the related weights of the task predictor (a Linear layer), one weight for each concept. In this way, it is visible the impact of each concept on each specific class. Then, we evaluate the impact of intervening on confounding concepts for class predictions [Koh et al., 2020]. We modify the dSprites task to highly correlate the colors with the shapes and therefore the task label. For instance, in 85% of samples with a positive label (presence of at least one square or heart), the objects are green, while in 85% of the samples with a negative label, the objects are red or blue. What if? (counterfactual actionability): drawing from previous works on counterfactuals, we compute: (i) the validity($\uparrow$[Wachter et al., 2017] by checking whether the model predicts the desired class labels based on the generated counterfactual; (ii) the proximity ($\downarrow$[Pawelczyk et al., 2020] which evaluates counterfactuals’ “reliability” as their similarity w.r.t. training samples; (iii) the time ($\downarrow$)  [Romashov et al., 2022] to generate a counterfactual for a set of samples.
@@
In addition, we propose the following metrics to assess the quality of counterfactuals: (i) the $\Delta$-Sparsity ($\downarrow$) which evaluates user’s “wasted efforts” by counting the number concepts changed [Guo et al., 2023] w.r.t. the minimal number of changes that would have generated a counterfactual according to the dataset:
@@
where $OptimalSparsity$ represents the mean Hamming distance between each concept vector in the test set and the closest concept vector of a drawn random class $y^{\prime}$, found using a “brute-force” approach and $Sparsity$ represents the mean Hamming distance between the predicted concepts $\hat{c}$ and the predicted concept counterfactuals $\hat{c^{\prime}}$; (ii) the “plausibility” [Wachter et al., 2017] as the Intersection over Union (IoU ($\uparrow$)) [Jaccard, 1912] between counterfactuals and ground-truth concept vectors; (iii) the variability ($\uparrow$) which evaluates counterfactuals’ “diversity”, as the cardinality ratio between the set of counterfactuals generated and the set of training concept vectors
@@
where $|c^{\prime}|$ represents the cardinality of the set composed by all $\hat{c^{\prime}}$, while $|c|$ represents the cardinality of the set composed by all $c$; (iv) finally, we measure the concept accuracy of generated interventions (Acc Int. ($\uparrow$)) w.r.t. ground-truth optimal interventions that evaluate the model’s ability to generate potential concept interventions. Following Espinosa Zarlenga et al. [2022], we inject noise on predicted concepts $\hat{c}$ to reduce task accuracy (randomly flip the value of some concepts), then we sample counterfactuals, conditioning on the ground-truth label $y$, and use them to fix mispredicted concept labels automatically.
@@
B.3 Baselines and Implementation details
Hyperparameters
To train our models and the baselines we selected the best hyperparameters according to our experiments. Table 3 shows the number of epochs, learning rate, and embedding size in the latent space, batch size for each dataset. They are shared among the baselines, and we took the best checkpoint out of the entire training for each model. In addition, Table 4 illustrates the parameters used to weight each term in the loss. However, all these details are in the released code.
Hyperparams dSprites MNIST add CUB
Epochs 75 150 150
Learning rate 0.005 0.005 0.005
Hidden size 128 128 128
Batch size 1024 1024 1024
@@ B.3 Baselines and Implementation details
Loss Term dSprites MNIST add CUB
CBM Concept 1.0 1.0 1.0
Task 0.1 0.1 0.1
CCHVAE Reconstruction 1.0 1.0 1.0
KL 0.5 0.5 0.5
VAECF Reconstruction 3.0 1.3 1.0
KL 1.0 1.0 2.0
Validity 15.0 15.0 15.0
VCNET Concept 1.0 1.0 1.0
Task 0.5 0.5 0.5
Reconstruction 0.5 0.5 0.5
KL 0.8 0.8 0.8
CF-CBM Concept 10.0 10.0 1.0
Task 0.7 1.0 0.1
Validity 0.3 0.2 0.02
$z$ KL 1.2 2.0 0.2
$z^{\prime}$ KL 1.2 2.0 0.2
Prior distance 1.0 1.7 0.2
Posterior distance 0.6 0.55 0.03
Table 4: Loss weights
Counterfactual CBM implementation details
@@
To improve the training process of our models we decided to relax some assumptions we made in Section 3.2. Employing a Bernoulli distribution to predict concepts and a Categorical distribution for the class prediction would make the training process more difficult, ruining the gradient in the backpropagation process. Therefore, we predict them directly with the output of the concept predictor and the task predictor applying a Sigmoid function and a Softamx on top of them, respectively. To optimise them, we choose to use a Binary Cross Entropy loss for the concept loss, and a Cross Entropy Loss for the task loss.
Code and licenses
@@
For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 [Paszke et al., 2019] (BSD license), and Sklearn 1.2 [Pedregosa et al., 2011] (BSD license). In addition, we used Matplotlib [Hunter, 2007] 3.7 (BSD license) to produce the plots shown in this paper. We will publicly release the code with all the details used to reproduce all the experiments under an MIT license.
Appendix C Further metrics and results
In addition to the result present in Section 5, Table 5 shows the task and concept performance of our model and the baselines. Moreover, Figure 8 illustrates the Acc Int. achieved by all models with different levels of noise injected at the concept level. It is visible how CF-CBM achieves stable Acc Int. results with different level of noise injected at the concept level. This is not the case for the other baselines. Finally, in contrast to Section 5, which shows an example of global feature importance for the dSprites dataset, Figure 9 represents the local feature importance of our model. Specifically, it illustrates how the model assesses the importance of features for predicting a specific class based on a single sample from the dataset.
dsprites MNIST add CUB
Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$)
CBM $99.6\pm 0.1$ $98.8\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.0\pm 0.0$ $85.3\pm 0.1$
BayCon $99.6\pm 0.0$ $98.7\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.1\pm 0.1$ $85.3\pm 0.1$
CCHVAE $99.5\pm 0.1$ $98.9\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.2\pm 0.1$ $85.4\pm 0.1$
VAECF $99.5\pm 0.0$ $98.8\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $92.9\pm 0.2$ $85.4\pm 0.1$
VCNET $99.7\pm 0.0$ $98.6\pm 0.1$ $98.2\pm 0.0$ $99.6\pm 0.0$ $89.6\pm 0.5$ $80.7\pm 0.1$
ACBM $99.7\pm 0.0$ $98.7\pm 0.1$ $97.0\pm 0.1$ $99.3\pm 0.0$ $91.4\pm 0.2$ $84.8\pm 0.2$

### v3

@@ C.2 Evaluation Metrics
In our analysis we use the following metrics. What? (prediction generalization): we compute the Area Under the Receiver Operating Characteristic Curve (Hand & Till, 2001) for concepts and tasks (ROC AUC ($\uparrow$). How? (intervention impact): we compute the impact of each concept on the downstream task, inspecting the task predictor $f$ (Yuksekgonul et al., 2022). For each class, we look at the related weights of the task predictor (a Linear layer), one weight for each concept. In this way, it is visible the impact of each concept on each specific class. Then, we evaluate the impact of intervening on confounding concepts for class predictions (Koh et al., 2020). We modify the dSprites task to highly correlate the colors with the shapes and therefore the task label. For instance, in 85% of samples with a positive label (presence of at least one square or heart), the objects are green, while in 85% of the samples with a negative label, the objects are red or blue. Why not? (counterfactual actionability): drawing from previous works on counterfactuals, we compute: (i) the validity($\uparrow$(Wachter et al., 2017) by checking whether the model predicts the desired class labels based on the generated counterfactual; (ii) the proximity ($\downarrow$(Pawelczyk et al., 2020) which evaluates counterfactuals’ “reliability” as their similarity w.r.t. training samples; (iii) the time ($\downarrow$)  (Romashov et al., 2022) to generate a counterfactual for a set of samples.
@@
In addition, we propose the following metrics to assess the quality of counterfactuals: (i) the $\Delta$-Sparsity ($\downarrow$) which evaluates user’s “wasted efforts” by counting the number concepts changed (Guo et al., 2023) w.r.t. the minimal number of changes that would have generated a counterfactual according to the dataset:
@@
where $OptimalSparsity$ represents the mean Hamming distance between each concept vector in the test set and the closest concept vector of a drawn random class $y^{\prime}$, found using a “brute-force" approach and $Sparsity$ represents the mean Hamming distance between the predicted concepts $\hat{c}$ and the predicted concept counterfactuals $\hat{c^{\prime}}$; (ii) the “plausibility” (Wachter et al., 2017) as the Intersection over Union (IoU ($\uparrow$)) (Jaccard, 1912) between counterfactuals and ground-truth concept vectors; (iii) the variability ($\uparrow$) which evaluates counterfactuals’ “diversity”, as the cardinality ratio between the set of counterfactuals generated and the set of training concept vectors
@@
where $|c^{\prime}|$ represents the cardinality of the set composed by all $\hat{c^{\prime}}$, while $|c|$ represents the cardinality of the set composed by all $c$; (iv) finally, we measure the concept accuracy of generated interventions (Acc Int. ($\uparrow$)) w.r.t. ground-truth optimal interventions that evaluate the model’s ability to generate potential concept interventions. Following Espinosa Zarlenga et al. (2022), we inject noise on predicted concepts $\hat{c}$ to reduce task accuracy (randomly flip the value of some concepts), then we sample counterfactuals, conditioning on the ground-truth label $y$, and use them to fix mispredicted concept labels automatically.
@@
C.3 Baselines and Implementation details
Hyperparameters To train our models and the baselines we selected the best hyperparameters according to our experiments. Table 5 shows the number of epochs, learning rate, and embedding size in the latent space, batch size for each dataset. They are shared among the baselines, and we took the best checkpoint out of the entire training for each model. In addition, Table 6 illustrates the parameters used to weight each term in the loss for all the methods. These hyperparameters were chosen through validation on a subset of the training to achieve the best possible trade-off across all metrics important for counterfactual evaluation, such as validity, proximity, and sparsity, for both our method and the baselines. While our method introduces more hyperparameters, it provides greater flexibility to specify the desired objective. As counterfactual metrics often conflict, creating inherent trade-offs (as shown in our experiments), there is no one-size-fits-all solution. Thus, optimizing the counterfactual generator based on the metric of interest is crucial, and our method facilitates this. Each hyperparameter in the loss has a specific role in steering the model toward a particular behavior:
  • $\lambda_{1}$ (task-related) and $\lambda_{2}$ (concept-related) are important in all configurations as it remains the main goal of the model.
  • $\lambda_{3}$ prioritizes validity, encouraging counterfactuals with predictions matching $y{\prime}$, though potentially compromising proximity or sparsity.
  • $\lambda_{4}$ and $\lambda_{5}$ emphasize generating more realistic counterfactuals, potentially reducing proximity, which may trade off against validity and sparsity.
  • $\lambda_{6}$ and $\lambda_{7}$ promotes sparsity by reducing the number of changes from the input, which may trade off against validity. However, we observed that $\lambda_{6}$ has minimal influence on the optimization process, with $\lambda_{7}$ being the key hyperparameter driving this behaviour.
This flexibility makes our model more versatile and practical compared to baselines, which have limited or partial support for these trade-offs (e.g., VAECF has the possibility to be flexible on validity). To show these behaviours we perform an ablation study on MNIST add that is shown in Table 7.
@@ C.3 Baselines and Implementation details
Hyperparams dSprites MNIST add CUB CIFAR10 SIIM Pneumothorax
Epochs 75 150 150 30 100
Learning rate 0.005 0.005 0.005 0.005 0.005
Hidden size 128 128 128 128 128
Batch size 1024 1024 1024 1024 1024
Table 6: Loss weights
@@
Loss Term dSprites MNIST add CUB CIFAR10 SIIM Pneumothorax
CBM Concept 1.0 1.0 1.0 1.0 1.0
Task 0.1 0.1 0.1 0.1 0.1
CCHVAE Reconstruction 1.0 1.0 1.0 1.0 1.0
KL 0.5 0.5 0.5 0.5 0.5
VAECF Reconstruction 3.0 1.3 1.0 1.3 1.3
KL 1.0 1.0 2.0 1.0 1.0
Validity 15.0 15.0 15.0 20.0 20.0
VCNET Concept 1.0 1.0 1.0 1.0 2.0
Task 0.5 0.5 0.5 0.5 0.5
Reconstruction 0.5 0.5 0.5 0.2
KL 0.8 0.8 0.8 0.8 0.8
CF-CBM Concept 10.0 10.0 1.0 10.0 10.0
Task 0.7 1.0 0.1 1.0 1.0
Validity 0.3 0.2 0.02 0.2 0.2
$z$ KL 1.2 2.0 0.2 2.0 2.0
$z^{\prime}$ KL 1.2 2.0 0.2 2.0 2.0
Prior distance 1.0 1.7 0.2 1.7 1.7
Posterior distance 0.6 0.55 0.03 0.4 0.4
Table 7: Performance comparison for different objective focuses with varying $\lambda$ values on the MNIST add dataset.
Objective Focus $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$ $\lambda_{7}$ Validity (%) Sparsity (%) Proximity (%)
Higher focus on validity 0.4 2 0.55 $99.1\pm 0.2$ $19.0\pm 0.4$ $4.3\pm 0.7$
Higher focus on proximity 0.2 4 0.55 $93.6\pm 1.6$ $14.2\pm 0.4$ $3.6\pm 0.0$
Higher focus on sparsity 0.2 2 0.70 $74.4\pm 0.01$ $6.1\pm 0.4$ $2.6\pm 0.0$
@@
Counterfactual CBM implementation details To improve the training process of our models we decided to relax some assumptions we made in Section 3.1. Employing a Bernoulli distribution to predict concepts and a Categorical distribution for the class prediction would make the training process more difficult, ruining the gradient in the backpropagation process. Therefore, we predict them directly with the output of the concept predictor and the task predictor applying a Sigmoid function and a Softamx on top of them, respectively. To optimise them, we choose to use a Binary Cross Entropy loss for the concept loss, and a Cross Entropy Loss for the task loss. All this is done in a jointly training to fully leverage the advantages outlined in Section 5.2. This means that all component are trained together. However, it is also possible to train the model in a post-hoc manner—first training the concept encoder and predictor, followed by the counterfactual generator, or performing a warm-up phase for the encoder and predictor to better initialize their weights before continuing with joint training.
@@
Code, licenses and hardware For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 (Paszke et al., 2019) (BSD license), PytorchLightning v2.1.2 (Apache Licence 2.0), Sklearn 1.2 (Pedregosa et al., 2011) (BSD license). In addition, we used Matplotlib (Hunter, 2007) 3.7 (BSD license) to produce the plots shown in this paper. The datasets we used are freely available on the web with licenses: dSprites (Apache 2.0) MNIST (CC 3.0 DEED) and CUB (MIT License). We will publicly release the code with all the details used to reproduce all the experiments under an MIT license. The experiments were performed on a device equipped with an M3 Max and 36GB of RAM, without the use of a GPU. Approximately 80 hours of computational time were utilized from the start of the project, whereas reproducing the experiments detailed here requires only 10 hours.
Appendix D Further metrics and results
In addition to the result present in Section 5, Table 8 shows the task and concept performance of our model and the baselines. Moreover, Figure 6 illustrates the Acc Int. achieved by all models with different levels of noise injected at the concept level. It is visible how CF-CBM achieves stable Acc Int. results with different level of noise injected at the concept level. This is not the case for the other baselines. Figure 7 shows the distribution of the CaCE score obtained in MNIST add and CUB in the post-hoc and joint fashions. It allows to directly see the distributions used to compute the metrics in Table 2. Finally, Figure 8 shows the time efficency of all the method in generating counterfactuals. Generation efficiency is a key feature for real-time applications. In this respect, CF-CBMs significantly outperform iterative models such as C-CHVAE and BayCon. These models quickly struggle in large search spaces involving a high number of concepts: generating a counterfactual takes more than a minute on dSprites (7 concepts) and up to a few hours on CUB (118 concepts). CF-CBMs instead scale well with the size of the data set by requiring less than a second on dSprites and a few seconds on CUB, on par with VAECF and VCNet. This efficiency enables a wider range of applications requiring a high number of concepts and quick feedbacks.
@@ Appendix D Further metrics and results
dsprites MNIST add CUB CIFAR10 SIIM Pneumothorax
Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$)
Black Box $99.9\pm 0.0$ $98.4\pm 0.0$ $94.4\pm 0.0$ $99.7\pm 0.0$ $95.1\pm 0.2$
CBM $99.6\pm 0.1$ $98.8\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.0\pm 0.0$ $85.3\pm 0.1$ $99.3\pm 0.0$ $99.1\pm 0.0$ $89.4\pm 0.6$ $99.6\pm 0.1$
CEM $99.8\pm 0.1$ $98.6\pm 0.1$ $98.4\pm 0.0$ $99.5\pm 0.0$ $93.3\pm 0.1$ $85.7\pm 0.1$ $99.8\pm 0.0$ $99.2\pm 0.0$ $94.2\pm 0.1$ $99.5\pm 0.0$
BayCon $99.6\pm 0.0$ $98.7\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.1\pm 0.1$ $85.3\pm 0.1$ $99.4\pm 0.0$ $99.1\pm 0.0$ $89.3\pm 0.8$ $99.6\pm 0.1$
CCHVAE $99.5\pm 0.1$ $98.9\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.2\pm 0.1$ $85.4\pm 0.1$ $99.3\pm 0.0$ $99.2\pm 0.0$ $89.6\pm 0.5$ $99.3\pm 0.1$
VAECF $99.5\pm 0.0$ $98.8\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $92.9\pm 0.2$ $85.4\pm 0.1$ $99.3\pm 0.0$ $99.2\pm 0.0$ $89.5\pm 0.8$ $99.6\pm 0.1$
VCNET $99.7\pm 0.0$ $98.6\pm 0.1$ $98.2\pm 0.0$ $99.6\pm 0.0$ $89.6\pm 0.5$ $80.7\pm 0.1$ $99.1\pm 0.1$ $97.4\pm 0.2$ $91.4\pm 0.7$ $99.2\pm 0.0$
CF-CBM $99.7\pm 0.0$ $98.7\pm 0.1$ $97.0\pm 0.1$ $99.3\pm 0.0$ $91.4\pm 0.2$ $84.8\pm 0.2$ $99.4\pm 0.0$ $96.6\pm 0.0$ $92.0\pm 0.5$ $98.2\pm 0.0$